Required information Use the following information for the Quick Study below. [The following information applies to the questions displayed below.] Brodrick Company expects to produce 21,600 units for the year ending December 31. A flexible budget for 21,600 units of production reflects sales of $583,200; variable costs of $64,800; and fixed costs of $144,000. QS 23-4 Flexible budget performance report LO P1 Assume that actual sales for the year are $703,600 (27,800 units), actual variable costs for the year are $113,200, and actual fixed costs for the year are $136,000. Prepare a flexible budget performance report for the year.
Answer #1
Flexible budget performance report
Flexible budget Actual results variance Favorable/unfavorable
sale 750600 703600 47000 unfavorable
variable expenses 83400 113200 29800 unfavorable
contribution margin 667200 590400 17200 unfavorable
fixed expenses 144000 136000 8000 Favorable
income from operation 523200 454400 9200 unfavorable
Working
sales (583200/21600)*27800 = 750600
variable expenses= 64800/21600*27800 = 83400

Please give a thumbs up if it is helpful & let me know if any doubt